Physics 2

1. An ideal gas is in a container under a pressure of 20.0 atm at a temperature of 27.0oC. If half of the gas is released from the container and the temperature rises to 50.0oC, what is the final gas pressure?

2. Find the speed rms of the Hydrogen (H2) molecules at a temperature of 27oC
Answer #1

Consider ideal gas equation

P V = n R T

P / n T = R / V = constant

Here P is pressure, n is number of mole , T is temperature
P1 / n1 T1 = P2 / n2 T2

Half gas released, n2 = 0.5n1

(20 atm) / (n)(273 + 27) = P2 / (0.5n)(273+50)

P2 = 10.77 atm = 10.77 x 101kPa

P2 = 1080 kPa

2.

Given temperature of hydrogen , T = 27 degree C

T = 300 K

mass of H2 molecule , m = 0.002 Kg/mol

rms speed = sqrt(3 * R * T/m)

rms speed = sqrt(3 * 8.314 * 300/.002)

rms speed = 1934 m/s